# FlexLLM: A System for Co-Serving Large Language Model Inference and Parameter-Efficient Finetuning

### v1

@@
Abstract
Parameter-efficient finetuning (PEFT) is a widely used technique to adapt large language models for different tasks. Service providers typically create separate systems for users to perform PEFT model finetuning and inference tasks. This is because existing systems cannot handle workloads that include a mix of inference and PEFT finetuning requests. As a result, shared GPU resources are underutilized, leading to inefficiencies. To address this problem, we present FlexLLM, the first system that can serve inference and parameter-efficient finetuning requests in the same iteration. Our system leverages the complementary nature of these two tasks and utilizes shared GPU resources to run them jointly, using a method called co-serving. To achieve this, FlexLLM introduces a novel token-level finetuning mechanism, which breaks down the finetuning computation of a sequence into smaller token-level computations and uses dependent parallelization and graph pruning, two static compilation optimizations, to minimize the memory overhead and latency for co-serving. Compared to existing systems, FlexLLM’s co-serving approach reduces the activation GPU memory overhead by up to 8$\times$, and the end-to-end GPU memory requirement of finetuning by up to 36% while maintaining a low inference latency and improving finetuning throughput. For example, under a heavy inference workload, FlexLLM can still preserve more than 80% of the peak finetuning throughput, whereas existing systems cannot make any progress with finetuning. The source code of FlexLLM is publicly available at https://github.com/flexflow/FlexFlow/.
@@ 1 Introduction
Recent advancements in generative large language models (LLMs) such as GPT [9, 22, 47], PaLM [14], and LLaMA [62] have shown strong capabilities of generating natural language texts across various application domains [72, 39, 60, 21]. The ongoing evolution of LLMs is accompanied by exponential increases in their model sizes, with some models encompassing more than 100 billion parameters [50, 53], making it infeasible to fully finetune an LLM for downstream tasks (i.e., training on task-specific datasets).
To reduce the computational requirement of LLM finetuning, recent work introduces parameter-efficient finetuning (PEFT), which updates a subset of trainable parameters from an LLM or introduces a small number of new parameters into the LLM while keeping the vast majority of the original LLM parameters frozen [28, 36, 48, 29, 26, 34]. As a result, PEFT only requires maintaining a small fraction of optimizer states for updating model parameters, dramatically reducing the memory requirement compared to full finetuning [17]. The recent progress of PEFT-based LLMs has proved that PEFT can achieve comparable performance to full finetuning while enabling fast adaption to new tasks [48].
Given the distinct needs of training and inference tasks, which are respectively optimized for throughput and latency, existing cluster schedulers generally deploy the two tasks separately [23, 11]. More importantly, current LLM runtime systems are designed to support either PEFT-based LLM finetuning task or inference tasks exclusively, but none can accommodate the concurrent serving of both task types [57, 10, 41]. As a result, researchers and practitioners have to allocate distinct GPU resources to LLM inference and finetuning tasks. This approach achieves suboptimal GPU utilization for both tasks, as shown in Figure 1. In particular, due to the autoregressive nature of LLMs, existing LLM inference systems use incremental decoding to perform inference computation where the system iteratively decodes one token using the input prompt and all previously generated tokens. Consequently, LLM inference is primarily bounded by memory access since the decoding of each token requires accessing all model parameters [32, 43]. On the other hand, finetuning an LLM can process all tokens of a request simultaneously (e.g., 4K tokens for LLaMA-2 [63]), making finetuning largely bottlenecked by compute resources such as the compute capability of tensor cores on modern GPUs.
Figure 1: Average GPU memory bandwidth and compute resources utilization for LLM finetuning and inference. The results are collected by serving and finetuning a LoRA-assisted LLaMA-2-70B model on four and eight NVIDIA A100 40GB GPUs. We use a maximum sequence length of 32K for inference and a sequence length of 4K for LoRA finetuning.
Based on a critical observation that LLM inference and finetuning tasks are complementary, this paper introduces FlexLLM, the first co-serving system for LLM inference and parameter-efficient finetuning [^1] For simplicity, we use the term finetuning to refer to parameter-efficient finetuning for the rest of the paper.. FlexLLM introduces a PEFT-as-a-service (PaaS) interface that unifies inference and finetuning tasks and jointly serves them on shared GPU resources. This co-serving approach allows FlexLLM to achieve the best of both worlds. Figure 2 shows the inference and finetuning performance achieved by FlexLLM through co-serving. By fusing the computation of inference and finetuning and allowing them to share model weights, FlexLLM outperforms the conventional approach of employing twice as many GPUs for resource isolation. Specifically, FlexLLM retains more than 80% of the peak throughput of finetuning tasks while simultaneously preserving the latency of inference tasks. Co-serving inference and finetuning requires addressing two key challenges. Next, we elaborate on these challenges and the main ideas FlexLLM uses to manage them.
Inference latency.
Most LLM inference tasks are online and dynamically fluctuating, necessitating a system that can rapidly adjust GPU resources to guarantee low latency. The inference of generative LLMs follows a token-by-token pattern, while a finetuning task processes an entire sequence in each iteration to maximize throughput. As a result, directly fusing the computation of an inference iteration with that of a finetuning iteration would significantly prolong the inference latency. For example, under the setting from Figure 1, an inference and a finetuning iteration take, respectively, 50ms and 1500ms.
To address this challenge, FlexLLM introduces a novel token-level finetuning mechanism, which decomposes the finetuning computation of a sequence into a finer granularity of multiple tokens (Section 6.1). This approach significantly reduces the per-iteration latency of finetuning for both forward and backward passes, and allows FlexLLM to finetune an arbitrary number of tokens for each iteration. FlexLLM’s hybrid token scheduler coordinates inference and finetuning tokens for each co-serving iteration to maximize resource utilization while attaining the service level objective (SLO) for inference requests (Section 6.2).
Memory overhead.
Within existing finetuning systems, the update process for a small group of model parameters necessitates the retention of all intermediate activations to facilitate the gradient computation of those parameters subject to training. The activations substantially contribute to the memory requirements during finetuning, making it more resource-intensive than inference. FlexLLM introduces fundamental techniques to minimize the memory requirement of PEFT. First, different from fully finetuning an LLM’s weights, PEFT methods generally keep all weights of the original LLM frozen and introduce multiple bypass networks with small amounts of trainable parameters. Based on this observation, FlexLLM makes all finetuning and inference tasks built on top of the same base LLM so that they can share model weights and intermediate activations. FlexLLM introduces dependent parallelization to discover an optimized parallelization strategy for a PEFT model by solving a constrained optimization problem (Section 5.1). Instead of preserving all intermediate activations during a forward pass for finetuning as existing systems, FlexLLM leverages a critical observation that most intermediate activations are used for computing the gradients for frozen parameters and performs graph pruning to eliminate data dependencies for frozen parameters (Section 5.2). These memory optimizations allow FlexLLM to reduce PEFT memory overhead by up to 8$\times$ compared to existing systems.
@@
Contributions.
This paper makes the following contributions:
  • We show that LLM inference and finetuning are complementary and present their first co-serving system.
  • We propose a novel mechanism to decompose finetuning computation into the token-level granularity, which allows fusing inference and finetuning computation.
  • We introduce dependent parallelization and graph pruning, two static compilation optimizations that reduce PEFT memory overhead by 8$\times$.
  • We build FlexLLM, a co-serving system that implements these techniques and vastly outperforms existing systems.
Figure 2: Co-serving LLAMA-2-70B inference and finetuning on four NVIDIA A100 40GB GPUs with different numbers of inference requests and finetuning tokens. Data points with a zero finetuning throughput report the latency when serving only inference requests, while other data points mix inference and finetuning workloads. The green line shows the peak finetuning throughput.
@@ 2.1 Parameter-Efficient Finetuning
Finetuning allows an LLM to adapt to specialized domains and tasks [18, 49, 9]. However, finetuning an entire LLM introduces significant computation overheads due to the very large model size. Parameter-efficient finetuning (PEFT) methods [20] are proposed to solve this issue. Drawing inspiration from the pronounced sensitivity of LLMs to prompts as highlighted in [54], a series of studies introduce trainable prompt embeddings prepended to the input text or attention components while preserving the original LLM parameters [40, 36, 34]. [52] and [28] propose adapter modules to introduce new task-specific parameters, which are inserted into the Transformer layers inside the LLM. LoRA [29] leverages the low-rank decomposition concept to construct trainable parameters inserted into the original LLM weights. (IA)${}^{3}$ [38] proposes to scale the pre-trained model weights of an LLM with a trainable vector. Of late, there has been a surge in the proposal of unified approaches that amalgamate various PEFT methods by leveraging human heuristics [26] or employing neural architecture search [76, 78, 42]. Existing PEFT approaches focus on optimizing model performance while minimizing trainable parameters. However, reducing the number of trainable parameters does not inherently imply a corresponding reduction in memory footprint.
Figure 3: Comparison of different sharing approaches for serving PEFT and inference simultaneously. For spatial sharing and our co-serving approaches, the height of rounded rectangles illustrates the splitting ratio of GPU resources (e.g., SMs or threads).
@@ 2.2 PEFT-based LLM Inference
LLM inference is increasingly essential in both research and industrial applications. Unfortunately, deploying high-quality LLMs without exhausting the available GPU resources can be difficult due to high computational demands and memory requirements. One of the key challenges in existing LLM inference systems is the memory-intensive auto-regressive decoding mechanism, since it requires accessing all of the model’s parameters to generate each token. Several techniques have been proposed to overcome the memory-bandwidth limitation, such as quantization [66], parallelization [70], batching [58], and kernel optimizations [16]. However, the memory-bandwidth problem still exists, especially in scenarios where both latency and response quality matter. At inference time, PEFT, which unlocks a resource-efficient way to finetune LLMs for specific tasks, also suffers from the same memory access challenges as backbone LLMs.
Recent works have introduced a diversity of systems for simultaneously serving multiple PEFT-based LLMs. PetS[77] is the first unified framework for multi-task Parameter-Efficient Transformers (PETs) serving, which decouples linear layers in the computation graph into task-agnostic shared operations and task-specific PET operations, then uses a high-performance inference engine an innovative operator scheduling strategy to improve concurrency and throughput. While PetS only supports encoder-only models, recent work also focuses on optimizations for generative inference models. PUNICA[10] is a system to serve multi-tenant LoRA in a shared GPU cluster. It proposes a new CUDA kernel called Segmented Gather Matrix-Vector Multiplication, which allows batching GPU operations for the concurrent execution of multiple, different LoRA models. S-LORA[57] is optimized on the kernel introduced by PUNICA. It proposes unified paging, heterogeneous batching, and a novel tensor parallel algorithm to serve thousands of LoRA components on a single machine. However, all these systems are also bounded by memory access like conventional LLM serving systems [32, 70].
@@ 3 Co-serving LLM Inference and PEFT
Previous studies have proposed different methods for scheduling heterogeneous workloads (in our case, inference and PEFT) on shared GPU resources. These methods can be classified into three categories: resource isolation, temporal sharing, and spatial sharing. This section will discuss the limitations of these approaches when it comes to serving LLM inference and PEFT, and highlight the critical differences between FlexLLM’s co-serving solution and previous work. Figure 3 illustrates the GPU execution timeline of different scheduling approaches for serving inference and finetuning.
@@
A straightforward approach to support both inference and finetuning is resource isolation. This involves partitioning the available GPUs into two isolated groups, each handling one type of request (Figure 3(a)). Resource isolation guarantees that throughput-sensitive finetuning requests do not interfere with latency-sensitive inference requests. When finetuning an LLM, users usually provide a dataset of requests to train the PEFT layers. All the requests in the dataset are submitted to FlexLLM at once, making it easier to build larger batches that can be consumed sequentially. On the other hand, each user submits inference requests independently and expects them to be processed within a pre-defined SLO, and the arrival time of each request is unknown. Unfortunately, resource isolation can introduce significant GPU under-utilization for workloads with fluctuating request arrival rates, because finetuning and inference requests cannot be batched together when the number of pending requests of a given type is insufficient to fill a batch on its own. With fewer GPUs available to parallelize the inference computations, it may not be possible to serve requests fast enough to meet their SLO (e.g., request $r_{3}$ in Figure 3(a)).
@@
Compared to resource isolation, a temporal sharing [67, 68] approach can increase GPU utilization by allowing inference and finetuning tasks to share GPU resources through time slices of different lengths depending on various parameters. This technique allows inference and finetuning requests to access all GPUs in the cluster in turns, as shown in Figure 3(b). Using all GPUs to parallelize computation (e.g., with data parallelism or tensor model parallelism) can help reduce the latency of each request, which in turn can help to meet inference SLO restrictions (e.g., inference request $r_{3}$ cannot be completed in Figure 3(a), but can be completed in Figure 3(b) by leveraging more resources). However, the arrangement of time slices can also be harmful because of the unpredictable arrival times of inference requests, which may have a stringent latency requirement that cannot be met (e.g., $r_{4}$ has been delayed in Figure 3(b) by pre-scheduled finetuning requests).
@@
An issue affecting both resource isolation and temporal sharing is that they do not take into account the different characteristics of compute-intensive finetuning tasks and memory-intensive inference tasks. This results in a significant waste of resources for both workloads. An alternative solution is spatial sharing, which allows running different tasks simultaneously, with each task occupying a fraction of GPU resources (i.e., streaming processors). For example, spatial sharing can be done on NVIDIA GPUs by writing multi-stream code or taking advantage of the Multi-Process Service (MPS) [4]. Further, on the latest Hopper and Ampere architectures, spatial sharing can take advantage of multi-instance GPU (MIG) [3] to adjust the partition configuration dynamically. However, the adjustment is not flexible or efficient [19, 69, 2] enough to accommodate the dynamic inference workload. Figures 3(c) and (d) show two separate cases where a larger fraction of GPU resources is assigned to finetuning or inference. While over-provisioning for inference tasks may seem to provide better SLO guarantees, it still fails in the event of unexpected and urgent requests like $r_{3}$. Additionally, over-provisioning wastes memory bandwidth resources and significantly slows down the finetuning process. Therefore, spatial sharing cannot be relied upon to run finetuning requests while also serving inference requests from online workloads with satisfactory performance [12, 75].
@@
As shown in Figure 3(e), FlexLLM takes a co-serving approach using a fine-grained scheduling mechanism to combine inference and finetuning requests adaptively. In each co-serving iteration, FlexLLM jointly serves inference and finetuning requests by (1) computing keys and values for inference requests in the pre-filling phase, (2) generating a new token for inference requests in the decoding phase, and (3) performing a forward and a backward pass for a subset of tokens from each finetuning request. FlexLLM fuses inference kernels with finetuning forward kernels to minimize kernel launch overhead and reduce accesses to GPU device memory to retrieve model weights. In each iteration, FlexLLM dynamically adjusts the allocation of inference and finetuning tokens, strategically balancing the need to meet the SLOs of inference requests with the goal of maximizing GPU utilization. This adaptive approach ensures that Collie adheres to the requirements of inference requests while optimizing the use of available GPU resources.
Compared to existing approaches, co-serving has two key advantages. First, FlexLLM achieves much higher GPU utilization than prior work. Both resource isolation and temporal sharing restrict each GPU to handle a single task type at any given time, leading to suboptimal utilization. This inefficiency stems from the fact that inference and finetuning tasks are constrained by GPUs’ compute and memory capabilities. Current spatial sharing mechanisms statically divide GPU resources for inference and finetuning, which cannot efficiently utilize GPUs for inference due to their inherently unpredictable nature. On the other hand, co-serving can be considered as a dynamic, iteration-level spatial sharing approach that adaptively adjusts resource allocation between inference and finetuning tasks in accordance with the inference and finetuning workloads.
Second, co-serving maximally retains the SLOs of inference tasks compared to temporal and spatial sharing. Temporal sharing violates stringent inference latency requirements when finetuning requests are pre-scheduled, and spatial sharing only allocates a subset of GPU resources for inference and therefore cannot effectively handle bursty inference requests. Co-serving addresses these issues by adaptively adjusting resource allocation based on the SLOs and the current workload of inference tasks.

### v2

@@
FlexLLM: A System for Co-Serving Large Language Model Inference and Parameter-Efficient Finetuning
Gabriele Oliaro     Xupeng Miao∗†     Xinhao Cheng     Vineeth Kada     Ruohan Gao     Yingyi Huang
Remi Delacourt     April Yang     Yingcheng Wang     Mengdi Wu     Colin Unger     Zhihao Jia
Carnegie Mellon University      Purdue University      Stanford University
footnotetext: Equal contribution.
Finetuning large language models (LLMs) is essential for task adaptation, yet today’s serving stacks isolate inference and finetuning on separate GPU clusters—wasting resources and under-utilizing hardware. We introduce FlexLLM, the first system to co-serve LLM inference and PEFT-based finetuning on shared GPUs by fusing computation at the token level. FlexLLM ’s static compilation optimizations—dependent parallelization and graph pruning significantly shrink activation memory, leading to end-to-end GPU memory savings by up to 80%. At runtime, a novel token-level finetuning mechanism paired with a hybrid token scheduler dynamically interleaves inference and training tokens within each co-serving iteration, meeting strict latency SLOs while maximizing utilization. In end-to-end benchmarks on LLaMA-3.1-8B, Qwen-2.5-14B, and Qwen-2.5-32B, FlexLLM sustains the inference SLO requirements up to 20 req/s, and improves finetuning throughput by 1.94.8×1.9-4.8$\times$1.9 - 4.8 × under heavy inference workloads and 2.56.8×2.5-6.8$\times$2.5 - 6.8 × under light loads, preserving over 76% of peak finetuning progress even at peak demand. The source code of FlexLLM is publicly available at https://github.com/flexflow/FlexFlow/.
@@ 1 Introduction
Recent advancements in large language models (LLMs) such as GPT [12, 24, 57], PaLM [17], and LLaMA [73] have shown strong capabilities of generating natural language texts across various application domains [88, 47, 68, 23]. Given the increasingly high cost of LLM pre-training, there is a trend of standardizing foundational pre-trained LLMs and sharing sub-models across multiple downstream tasks [34, 86] through finetuning (i.e., continuous training on small task-specific datasets). For example, a series of parameter-efficient finetuning (PEFT) techniques [34, 43, 59, 35, 32, 41] have been proposed to update a subset of trainable parameters from an LLM or introduce a small number of new parameters into the LLM while keeping the vast majority of the original LLM parameters frozen. The base LLM sharing paradigm provides opportunities to multiplex resources and improve cluster utilization (e.g., Google [11], Microsoft [35], and Tencent [56]). Many recent efforts [97, 80, 66, 13, 91, 36] inspired by these opportunities to build multi-task serving systems, handling inference requests for multiple fine-tuned LLMs simultaneously in a single GPU cluster.
Another undeniable trend is the increasing importance of LLM finetuning [^1] For simplicity, we use the term finetuning to refer to PEFT for the rest of the paper.. Almost all LLM serving companies, including OpenAI GPT-4o [58], Google Gemini [28], and Databricks Mosaic AI [19], now offer commercial finetuning APIs to facilitate the creation of custom models. This trend is not only due to the widespread adoption of LLMs but also driven by the growing demand for task-specific model customization [49], data access control [69], and the need for continuous and frequent model updating [26]. Unlike inference tasks with strict latency requirements (e.g., microseconds to milliseconds per token), finetuning tasks usually allow for greater latency tolerance with coarser time budget granularity (e.g., minutes to hours). As such, the current practice involves operating these two tasks on dedicated clusters [78, 27, 14], which exclusively use hardware resources through specialized systems optimized for latency and throughput, respectively.
However, real-world LLM inference workloads usually exhibit fluctuating and bursty request arrival patterns [77]. Serving inference requests must meet latency-based service level objectives (SLOs), which may vary from stringent to relaxed depending on the application scenario. To avoid SLO violations, service providers must either over-provision GPU resources or dynamically adjust resource allocation. The former strategy can lead to significant computational waste, while the latter introduces substantial overhead due to the need for frequent redeployments and process switching [10, 54]. In this paper, we explore a key research question: Can we design a system that simultaneously serves both inference and finetuning tasks within a shared GPU cluster, effectively accommodating the time-varying inference computational demands?
We introduce FlexLLM, the first co-serving system for LLM inference and finetuning. FlexLLM introduces a PEFT-as-a-service (PaaS) interface, which unifies inference and finetuning tasks, enabling their joint execution on shared GPU resources. The key insight behind FlexLLM is that inference and finetuning tasks, when using the same base LLMs, can be merged rather than executed separately.
This co-serving approach presents multiple advantages. First, since the base LLM is generally frozen during finetuning, the base model parameters can be stored without redundancy, therefore reducing the memory footprint. Second, when faced with bursty inference demands, FlexLLM can adaptively reassign GPU resources by temporarily reducing the proportion of finetuning tasks, which inherently possess higher latency tolerance. Once the peak inference workload has passed, the finetuning tasks can reclaim resources, thereby optimizing overall throughput and maximizing GPU utilization without compromising inference SLOs. This dynamic strategy provides a novel avenue for cloud LLM service providers to improve operational cost-effectiveness and service capabilities.
Beyond supervised PEFT, a growing class of reinforcement-learning-based fine-tuning algorithms – most notably Group Relative Policy Optimization (GRPO) used in DeepSeek-R1 [31], and multi-objective variants thereof [44] – require the model to generate complete sequences on-the-fly in order to compute policy gradients or reward signals. In each RLHF [65] or GRPO update, the LLM must sample continuations (i.e., perform inference), score them against a learned or hand-crafted reward, and then backpropagate through those same tokens. Because inference and policy-gradient steps are interleaved at the token level, splitting inference and training onto separate GPUs incurs costly data movement and context-switching. This tight coupling further motivates FlexLLM ’s token-level co-serving, which fuses generation and gradient updates in a single pass while still meeting stringent inference SLOs.
Challenges.
Co-serving inference and finetuning tasks within a single system, as envisioned by FlexLLM, presents significant challenges due to their differing workload characteristics. FlexLLM must address three key challenges.
First, how can inference and finetuning tasks be co-located while minimizing the memory footprint? Traditional multi-tenancy ML scheduling systems share resources across tasks and models [81], but suffer from the high memory demands associated with LLMs. Furthermore, existing finetuning systems usually retain all intermediate activations similar to training systems to enable gradient computation during backpropagation, which substantially increases their resource consumption compared to pure inference tasks. This discrepancy makes it challenging to manage system resources between finetuning and inference tasks.
To address this challenge, FlexLLM co-locates finetuning tasks with inference tasks by enforcing them to share the base LLM parameters and intermediate activation buffers. In scenarios where the base LLM has already been deployed in a distributed system, FlexLLM uses dependent parallelization to optimize the placement of small, trainable bypass networks (e.g., LoRA adapters [35]) introduced during finetuning. This is achieved by solving a constrained optimization problem detailed in §5.1.
Unlike existing systems that preserve all intermediate activations throughout a forward pass for finetuning, FlexLLM leverages the observation that most intermediate activations are only necessary for computing gradients of frozen parameters. By implementing graph pruning, FlexLLM eliminates unnecessary data dependencies associated with these frozen parameters (see §$\S$ 5.2). These memory optimization strategies allow FlexLLM to reduce the memory overhead associated for finetuning by up to 8$\times$ compared to existing systems.
Next, how can we run inference and finetuning tasks simultaneously while minimizing the interference with inference latency and SLO attainment? Previous work on traditional ML workloads [62] has shown that co-locating inference with training can significantly degrade inference performance. Unlike conventional training workloads that require coarse-grained temporal or spatial resource sharing [15], LLM finetuning provides a unique opportunity to batch base LLM computation and parameter access across requests from different finetuned variants [80]. However, due to differences in execution logic—auto-regresive decoding in inference versus backpropagation in finetuning—merging the two types of tasks remains difficult.
To overcome this challenge, FlexLLM introduces a novel token-level finetuning mechanism, which decomposes the finetuning computation of a sequence into smaller, token-level units (§$\S$6.1). During the forward pass, finetuning tokens follow the same operator execution order as decoding tokens in inference requests. These tokens are then processed using fused GPU kernels. For the backward pass, FlexLLM launches separate GPU streams for finetuning tokens and adopts a layer-wise execution strategy to minimize memory usage. This design significantly reduces the performance impact of finetuning tokens on inference latency, enabling efficient co-serving of inference and finetuning tasks on shared hardware.
Finally, how can we handle fluctuating inference workloads while preserving their latency SLO? Given the dynamic nature of inference request arrival patterns, provisioning for peak demand often results in significant resource under-utilization. FlexLLM’s fine-grained, token-level execution strategy enables quick adaptation to load fluctuations. Inspired by prior work on co-locating multiple cloud workloads [25], FlexLLM prioritizes inference tokens to meet strict latency constraints, while inserting finetunign tokens in a best-effort manner.
To achieve this, FlexLLM introduces a hybrid token scheduler, which coordinates the scheduling of inference and finetuning tokens at each co-serving iteration. This scheduler aims to maximize overall GPU utilization while guaranteeing that inference tasks consistently meet their SLOs (Section 6.2).
@@ 2.1 Parameter-Efficient Finetuning
Finetuning enables LLMs to adapt to specific domains and downstream tasks [20, 61, 12]. However, full finetuning of an LLM introduces significant computational and memory overheads due to the model’s large parameter count. To address this issue, a class of parameter-efficient finetuning (PEFT) methods [22] has been recently proposed. Inspired by the sensitivity of LLMs to input prompts [63], a series of studies introduce trainable prompt embeddings—either prepended to the input sequence or incorporated into attention components—while keeping the original LLM parameters frozen [48, 43, 41]. Recent work also proposed a variety of adapter modules to insert lightweight task-specific modules into the Transformer layers. For example, LoRA [35] leverages the low-rank decomposition concept to construct trainable parameters inserted into the original LLM weights. (IA)3 [46] proposes to scale the pre-trained model weights of an LLM with a trainable vector. Recently, there has been a surge in the proposal of unified approaches that amalgamate various PEFT methods by leveraging human heuristics [32] or employing neural architecture search [96, 98, 51]. Existing PEFT approaches focus on optimizing model performance while minimizing trainable parameters. However, reducing the number of trainable parameters does not inherently imply a corresponding reduction in memory footprint.
Figure 1: Comparing different resource sharing approaches for serving finetuning and inference. For spatial sharing and co-serving, the height of rounded rectangles illustrates the splitting ratio of GPU resources (e.g., streaming multi-processors).
@@ 2.2 PEFT-based LLM Inference
LLM inference is increasingly essential in both research and industrial applications. Recent work has introduced a diversity of systems for simultaneously serving multiple PEFT-based LLMs. PetS [97] is the first unified framework for multi-task parameter-efficient Transformers serving, which decouples linear layers in the computation graph into task-agnostic shared operations and task-specific PET operations, then optimizes the operator scheduling strategy to improve concurrency and throughput. While PetS only supports early encoder-only models, recent work also focuses on optimizations for generative decoding models. Punica [13] is a system to serve multi-tenant LoRA in a shared GPU cluster. It proposes a new CUDA kernel called Segmented Gather Matrix-Vector Multiplication, which allows batching GPU operations for the concurrent execution of multiple, different LoRA models. S-LoRA [66] is optimized on the kernel introduced by Punica. It proposes unified paging, heterogeneous batching, and a novel tensor parallel algorithm to serve thousands of LoRA components on a single machine. dLoRA [80] further improves efficiency by dynamically adjusting the LoRA batching plans under imbalanced workloads.
2.3 Reinforcement-Learning Based Finetuning
While PEFT methods such as LoRA and adapters optimize a small set of task parameters via supervised losses, recent work has demonstrated that pure reinforcement learning (RL) can substantially enhance LLM reasoning and alignment. In RLHF (e.g., [67]), the model samples candidate completions, a human or reward model ranks them, and policy gradients adjust the generator to favor preferred outputs. More recently, Group Relative Policy Optimization (GRPO) forgoes a separate value network by scoring groups of sampled continuations to estimate baselines, reducing critic overhead [31]. Extensions such as multi-objective GRPO further handle safety and helpfulness rewards within a single RL loop [44], and open-source systems like DAPO illustrate scaling these methods to 32B-parameter models [85].
Crucially, all of these RL-based pipelines require on-the-fly LLM generations as part of the reward calculation, interleaving inference kernels with backward passes at sub-sequence granularity. This pattern conflicts with the traditional separation of “inference-only” versus “training-only” clusters, leading to GPU underutilization or high-latency context switches. FlexLLM ’s token-level fusion directly addresses this challenge by co-scheduling generation and gradient updates within a single GPU pipeline, while still meeting strict inference SLOs.
@@ 3 Co-serving LLM Inference and PEFT
Previous studies have proposed different methods for scheduling heterogeneous workloads (in our case, inference and PEFT) on shared GPU resources. These methods can be classified into three categories: resource isolation, temporal sharing, and spatial sharing. This section discusses the limitations inherent to these conventional approaches when applied to co-serving LLM inference and PEFT. It also highlights the fundamental distinctions between FlexLLM’s co-serving solution and previous work. Figure 1 depicts the GPU execution timelines associated with different scheduling approaches for managing inference and finetuning tasks.
@@
A common approach to support both inference and finetuning tasks is resource isolation, which partitions the available GPUs into two distinct groups, each dedicated to handling one type of workload (see Figure 1(a)). This approach ensures that throughput-intensive finetuning tasks do not interfere with latency-critical inference tasks. When finetuning an LLM, a dataset of requests is provided to train the PEFT layers, with all requests submitted to FlexLLM simultaneously, making it easier to build larger batches that can be processed sequentially. In contrast, inference requests are submitted independently by users who expect timely processing within predetermined SLOs, with the arrival times of these requests unpredictable.
Resource isolation can introduce significant GPU under-utilization, particularly in scenarios with variable request arrival rates. This is because finetuning and inference requests cannot be batched together if the number of pending requests of either type is insufficient to fully utilize a batch. With fewer GPUs available to parallelize inference computations, it may not be possible to serve requests fast enough to meet their SLO (e.g., request $r_{3}$ in Figure 1(a)).
@@
Compared to resource isolation, temporal sharing [81, 82] improves GPU utilization by allowing different tasks to share GPU resources through time slices of different lengths depending on various parameters. This method allows both inference and finetuning requests to access all GPUs within the cluster in turns, as shown in Figure 1(b). By using all GPUs to parallelize computation (e.g., with data parallelism or tensor model parallelism), the latency for each request can be significantly reduced, thus aiding in meeting inference SLOs. For example, while inference request $r_{3}$ fails to complete in the scenario illustrated in Figure 1(a) due to resource limitations, it successfully executes in Figure 1(b) by leveraging more resources. However, the arrangement of time slices can also be harmful because of the unpredictable arrival times of inference requests, which may have a stringent latency requirement that cannot be met (e.g., $r_{4}$ has been delayed in Figure 1(b) by pre-scheduled finetuning requests).
@@
Spatial sharing offers an alternative by enabling simultaneous execution of different tasks at the kernel level, where each task utilizes a specific portion of the GPU resources (i.e., streaming multi-processors) This can be implemented on NVIDIA GPUs through multi-stream programming or by utilizing the Multi-Process Service (MPS) [4]. Furthermore, on the latest Hopper and Ampere architectures, spatial sharing benefits from the Multi-Instance GPU (MIG) [3], which allows for dynamic adjustments of GPU partition configurations. However, these adjustments often lack the flexibility and efficiency required to handle dynamic inference workloads effectively [21, 83, 2]. Figures 1(c) and (d) show two scenarios where a substantial portion of GPU resources is allocated to either finetuning or inference tasks. While over-provisioning for inference tasks may appear to provide better SLO guarantees, it still falls short during peak demand periods such as the unexpected urgent request $r_{3}$. Additionally, over-provisioning wastes resources and significantly slows down the finetuning process. Therefore, relying solely on spatial sharing to manage finetuning and inference tasks often results in suboptimal performance  [15, 92].
@@
As shown in Figure 1(e), FlexLLM employs a co-serving approach that uses a fine-grained scheduling mechanism to adaptively manage both inference and finetuning requests. During each co-serving iteration, FlexLLM jointly serves inference and finetuning requests by (1) computing keys and values for inference requests in the pre-filling phase, (2) generating a new token for inference requests in the decoding phase, and (3) performing forward and/or backward passes for selected tokens from finetuning requests.
Unlike temporal or spatial sharing approaches, which schedule GPU kernels of existing systems independently, FlexLLM integrates inference and finetuning kernels. This integration reduces the overhead associated with kernel launches and minimizes accesses to GPU device memory to retrive model weights. In each iteration, FlexLLM dynamically adjusts the allocation of inference and finetuning tokens, strategically balancing the need to meet the SLOs of inference requests with the goal of maximizing GPU utilization. This adaptive approach ensures that FlexLLM adheres to the requirements of inference requests while optimizing the use of available GPU resources.